OpenStudy (anonymous):

What is the solution to the equation? a + (–0.4) = –3 a =

OpenStudy (rushwr):

-2.6

OpenStudy (anonymous):

How can i get someone to help me with my question!

OpenStudy (rushwr):

\[a + (-0.4) = a- 0.4 \] \[a - 0.4 = -3\] \[a = - 2.6\]

OpenStudy (anonymous):

thank you @Rushwr

OpenStudy (anonymous):

oh, ok

OpenStudy (rushwr):

np :)